OpenStudy (gigi1103):

3x + 2y = 2 -2x + y = 8 Solve the system of equations. A) (-2,4) B) (4,-2) C) (-4,-2) D) (5 4 , -1 2 )

OpenStudy (gigi1103):

d(5/4 -1/2

OpenStudy (xaze):

Does D. contain fractions?

OpenStudy (gigi1103):

yes

OpenStudy (xaze):

This question falls into the same category that we did before. Let's start with A (-2,4) Can you try plugging in -2 for x and 4 for y?

OpenStudy (gigi1103):

ok 3(-2) +2 (4)=2

OpenStudy (xaze):

Now try solving. What is 3*-2 and 2*4?

OpenStudy (gigi1103):

3*-2= -6

OpenStudy (gigi1103):

2*4 = 8

OpenStudy (xaze):

Good. Now, does \[-6 + 8 = 2?\]

OpenStudy (gigi1103):

yes

OpenStudy (xaze):

Very nice. Try plugging in -2 for x and 4 for y, in the next part: \[-2x + y = 8\] and solve

OpenStudy (gigi1103):

-2(-2) +4=8

OpenStudy (gigi1103):

ok -2*-2=4 4+4 = 8

OpenStudy (gigi1103):

so its right can you check like one or two more with me

OpenStudy (santana):

this is easy

OpenStudy (gigi1103):

yeah i just want to be correct :T

OpenStudy (santana):

you have solve for (y) by getting it by itself and then you do the same for the other equation and you get an equation is this based on functions